Please can someone help Me this question
Please show the formula and the answe.
media%2Fd9d%2Fd9d5e2f0-ec52-460c-b37c-a0
Park Ridge Homecare Statement of Income Year Ended December 31, 2015 (in Thousands) Net Service Revenues Cost of Service Revenues 2666 1944 Gross Profit 722 General and Administrative Expenses Depreciation 649 15 Total Operating Expenses Operating Income 684 58 Interest Income Interest Expense 86 19 Total Interest Income, net Income Before Income Taxes 67 125 26 Income Tax Expense Net Income (Excess of Revenues over Expenses) Calculate the following Ratios below and compare with the ratio in 2016 and the Industry average. Insert Excel formula in the cells. Make sure you put the parenthesis in the formula to specify the order of operations to be done by the Excel formula a. Total (profit) Margin in 2015 Total (profit) Margin in 2016 Industry Average 40% 5.9% b. Operating Margin in 2015 3.2% 5.6% Operating Margin in 2016 Industry Average c. Return on Assets (ROA) in 2015 Industry Average d. Return on Equity (ROE) in 2015 Industry Average 13.4% 12.7% Return on Assets (ROA) in 2016 34.1% 36.5% Return on Equity (ROE) in 2016 e. Current Ratio in 2015 Current Ratio in 2016 Industry Average 1.9 days 24.7 days 22.6 days f. Days Cash on Hand in 2015 Days Cash on Hand in 2016 Industry Average
h. Debt Ratio in 2015 Debt Ratio in 2016 Industry Average 60.8% 63.9% i. Debt to Capitalization Ratio in 2015 Debt to Capitalization Ratio in 2016 Industry Average 24.4% 31.2% j. Times Interest Earned (TIE) Ratio in 2015 Times Interest Earned (TIE) Ratio in 2016 Industry Average times 13.3 times 10.4 times times 50.4 times 63.1 times k. Fixed Asset Turnover Ratio in 2015 Fixed Asset Turnover Ratio in 2016 Industry Average times 3.3 times 3.4 times I. Total Asset Turnover Ratio in 2015 Total Asset Turnover Ratio in 2016 Industry Average days 66.5 days 37.1 days m. Days in Accounts Receivable in 2015 Days in Accounts Receivable in 2016 Industry Average Interpret at least three Ratios by comparing the ratios in 2015, 2016 and the industry average. Is Park Ridge Homecare profitable? What areas should management be concerned about? Explain your answer.
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Profit margin in 2015= Net Income/ Net Service revenues

=99/ 2666

=3.71%

The profit margin is much below the margin in 2016. This is a good sign since the margin has increased showing higher profitability. It is also much lower than the industry average and the firm needs to analyze its cost and revenue structure.

2: Operating margin= Operating Income/ Net Service revenues

= 58/2666

=2.18%

The Operating margin is much below the margin in 2016. This is a good sign since the margin has increased showing higher profitability. It is also much lower than the industry average and the firm needs to analyze its operating costs.

3: ROA= Net Income/ Totalassets

= 99/879 = 11.26%

The ROA is much below the ROA in 2016. This is a good sign since the ratio has increased showing greater efficiency of assets used. It is much lower than the industry average which means the firm was less efficient in 2015, but in 2016 that has been covered up.

4: ROE= Net Income/ Equity

= 99/315

=31.43%